OpenStudy (anonymous):

MEDAL FOR CORRECT ANSWER PLZ HELP! If m<6=125degrees, find m<5 A)35deg B)55deg. C)65deg. D)125deg.

OpenStudy (danjs):

do you recall, the sum of the angles for a straight line are 180? so angle 6 plus angle 3 = 180

OpenStudy (danjs):

And the sum of the angles of a triangle are 180 angles 1+2+3 = 180

OpenStudy (danjs):

Angle 3 = 180 - 125 = 55 degrees

OpenStudy (anonymous):

I think i get it meow thank you lovely!

OpenStudy (danjs):



OpenStudy (anonymous):

I also got 35 after you explained the whole thing thank you so muchh! :)

OpenStudy (danjs):

welcome, u still have more to do?

OpenStudy (anonymous):

Yeah, I forgot how to do a few of these for the test coming up. For example Which is the equation of a line that is parallel to 2x+5y=4 and goes through the point 5,-4

OpenStudy (danjs):

ah, can you paste that in a new question thread, ill jump over to it

OpenStudy (anonymous):

okie dokie!